An investor is considering the purchase of Gryphon stock, which has returns given in the table below.

Scenario Probability Rate of Return
Recession 0.27 2%
Normal 0.58 9%
Boom 0.15 14%

Calculate the expected return and standard deviation of Gryphon. Round your answers to 2 decimal places.

Enter your answers below.

E(r) =    
Correct response: 7.86 %

Std. Dev. =   
Correct response: 3.96±0.01 %

The investor decides to diversify by investing $8,000 in Gryphon stock and $7,000 in Royal stock which has an expected return of 8.5% and a standard deviation of 9.5%. The correlation coefficient for the two stocks' returns is 0.2. Calculate the expected return and standard deviation of the portfolio. Round your answers to 2 decimal places. Use the correct answers from the previous question.

Enter your answers below.

E(rp) =  %

Std. Dev =  %
Answer #1

1-a). Expected Return = \sum_{i=1}^{n} [Probability(i) * Return(i)]

= [0.27 * 2%] + [0.58 * 9%] + [0.15 * 14%] = 0.54% + 5.22% + 2.10% = 7.86%

1-b). Standard Deviation = [\sum_{i=1}^{n}{Probability(i) * (Expected Return - Return(i))2}]1/2

= [{0.27 * (7.86% - 2%)2} + {0.58 * (7.86% - 9%)2} + {0.15 * (7.86% - 14%)2}]1/2

= [9.27%2 + 0.75%2 + 5.65%2]1/2

= [15.68%2]1/2 = 3.96%

2-a). Expected Return = \sum_{i=1}^{n} [Probability(i) * Return(i)]

= [(8,000/15,000) * 7.86%] + [(7,000/15,000) * 8.5%] = 4.19% + 3.97% = 8.16%

2-b). Standard Deviation = [{w(G)2 * S.D.(G)2} + {w(R)2 * S.D.(R)2} + {2 * w(G) * w(R) * S.D.(G) * S.D.(R) * Correlation(G,R)}]1/2

= [{(8,000/15,000)2 * (3.96%)2} + {(7,000/15,000)2 * (9.5%)2} + {2 * (8,000/15,000) * (7,000/15,000) * 3.96% * 9.5% * 0.2}]1/2

= [4.46%2 + 19.65%2 + 3.75%2]1/2

= [27.86%2]1/2 = 5.28%